# Enhancing Job Salary Prediction with Disentangled Composition Effect Modeling: A Neural Prototyping Approach

### v1

@@ 4.2.2. Skill-wise External Information Fusion
(3) $\begin{split}E^{\text{s}}_{h}&=\text{pool}\left(\{\alpha^{\text{lv}}_{i}\phi(x_{i})|x_{i}\in S_{h}\}\right).\end{split}$
@@
Figure 3. Illustrative examples of co-occurrence graphs. Low-weight edges are eliminated to ease presentation.
4.2.3. Co-occurrence Graph Density Regularization
To enhance explainability and robustness, extracted skill subsets should reflect meaningful real-world interactions. Yet, focusing solely on prediction performance could generate subsets cluttered with irrelevant skills, reducing semantic clarity. To address this, we introduce a regularization approach aimed at refining these subsets to meet key intuitive criteria: (i) Semantic coherence within subsets, (ii) Distinct semantic boundaries between different subsets, and (iii) Avoidance of trivial subset selections (neither fully inclusive nor empty sets).
Observing that skills frequently appeared together in job postings likely share semantic connections (e.g., “Machine Learning”, “Algorithms”, and “Python” often cluster in AI roles), we leverage global semantic interactions inferred from such co-occurrence patterns. Direct co-occurrence analysis, however, is computationally expensive and hampered by data sparsity. Therefore, we model these relationships using a skill co-occurrence graph $G_{skill}=\langle U,L\rangle$, drawing from skill pair associations across the dataset, where $U$ is the universal set of skills and $L$ represents the links indicating co-occurrence frequencies between skill pairs. As exemplified in Figure 3, each graph node corresponds to a skill, and a link $l_{i,j}$ between nodes $x_{i}$ and $x_{j}$ signifies their connectivity, quantified by their co-occurrence frequency $w_{l_{i,j}}$. To quantify semantic connections within subsets, we design a graph density-based (Khuller and Saha, 2009) regularization, focusing on the subgraph $G_{h}=\langle S_{h},L_{G_{h}}\rangle$ for each subset $S_{h}$. The subgraph’s density, calculated as $\frac{\sum_{l\in L_{\mathrm{G}_{h}}}w_{l}}{|L_{{\mathrm{G}_{h}}|}}$, gauges the semantic cohesiveness of the subset. By maximizing this density, we encourage subsets that are semantically well-connected, while minimizing trivial or overly inclusive selections. Our graph regularization of the extracted subsets $\{S_{h}\}_{h=1}^{H}$ is thus defined as:
(4) $\mathcal{L}_{con}=-\log\frac{1}{H}\sum_{h=1}^{H}\frac{\sum_{l\in L_{\mathrm{G}_{h}}}w_{l}}{|L_{\mathrm{G}_{h}}|}.$
@@ 4.2.3. Co-occurrence Graph Density Regularization
4.3. Prototypical Set Learning
Based on extracted subsets, LGDESetNet matches the input with learnable prototypes representing global semantics and accordingly models the output as an aggregation of these global contributions.
4.3.1. Set-oriented Prototypical Part Layer
@@
We assess the alignment between the input set and each prototype (i.e., global semantics) by comparing the similarity of extracted subsets to the prototypes. Notably, the discrete set representation brings the curse of dimensionality, complicating set similarity estimation. For instance, “C” and “C++” are treated as distinct skills despite their semantic closeness, potentially hindering model convergence and accuracy. To address this, we project prototype $P_{i}$ and subset $S_{h}$ into semantic representation vectors with the same skill embedding table as the input set, denoted as $E^{\text{p}}_{i}=\text{pool}\left(\{\gamma^{\text{s}}_{i}\gamma^{\text{lv}}_{i}\phi(x_{i})|x_{i}\in U\}\right)$ and $E^{\text{s}}_{h}=\text{pool}\left(\{\alpha^{\text{lv}}_{i}\phi(x_{i})|x_{i}\in S_{h}\}\right)$. Considering the distribution shift between the whole set and single-view subsets, we introduce a transformation layer $\mathcal{T}$ to further project the subsets and prototypes into a subset representation space as $Z^{\text{p}}_{i}=\mathcal{T}\left(E^{\text{p}}_{i}\right)$ and $Z^{\text{s}}_{h}=\mathcal{T}\left(E^{\text{s}}_{h}\right)$. In this way, we better capture semantic-level associations while ensuring the similarity can be measured by aligned representation of prototypes and subsets.
@@
(5) $sim(S_{h},P_{i})=\text{log}\left(\frac{\|Z^{\text{s}}_{h}-Z^{\text{p}}_{i}\|_{2}^{2}+1}{\|Z^{\text{s}}_{h}-Z^{\text{p}}_{i}\|_{2}^{2}+\epsilon}\right).$
@@
(6) $y=\sum_{i=1}^{M}\gamma^{\text{sal}}_{i}\cdot\text{softmax}(\sum_{h=1}^{H}sim(S_{h},P_{i})).$
4.3.2. Set-wise External Information Fusion
In practice, each skill set is enriched with set-wise external information, which is also useful in determining salary outcomes. For example, job salaries are highly correlated with the work experience and the pricing level of the base city. However, directly fusing such information into deep and wide embeddings (Cheng et al., 2016) could obscure the model’s explainability. Therefore, we separate the set-wise external information from skill embeddings, acknowledging that the salary effects of global prototypical skill sets may vary across different job scenarios. Specifically, we model this prototype salary effect dynamic as a function of the set-wise external features $C=\{c_{1},c_{2},...,c_{l}\}$. Drawing inspiration from DeepFM (Guo et al., 2017), we utilize a factorization machine encoder for low-order feature interactions and an MLP encoder for high-order interactions. By aggregating the embeddings from both encoders, we then employ another MLP to estimate the salary effects of each prototype, ensuring a dynamic and contextual assessment of skill sets’ impact on salary predictions:
(7) $\begin{split}\gamma^{\text{sal}}&=\text{MLP}\left(\text{FM}(C)+\text{MLP}(c_{1}|c_{2}|...|c_{l})\right).\\ \text{FM}(C)&=w_{0}+\sum_{i=1}^{l}w_{i}c_{i}+\sum_{i=1}^{l}\sum_{j=i+1}^{l}c_{i}\odot c_{j},\end{split}$
@@ 4.3.2. Set-wise External Information Fusion
4.3.3. Prototype Learning Regularization
To enhance explainability, we introduce regularization methods for our prototypical set learning procedure, whose basic ideas are widely adopted in prototype learning for other data structures (Zhang et al., 2022a; Ming et al., 2019). First, since prototypes are desired to be representative exemplars for the subsets, we introduce a representation regularization term to encourage each subset to be close to at least one prototype and encourage each prototype to be close to at least one extracted subset.
(8) $\mathcal{L}_{rep}=\frac{1}{H}\sum\limits_{h=1}^{H}\min_{i\in[1,M]}\|Z^{\text{s}}_{h}-Z^{\text{p}}_{i}\|_{2}^{2}+\frac{1}{M}\sum_{i=1}^{M}\min_{h\in[1,H]}\|Z^{\text{s}}_{h}-Z^{\text{p}}_{i}\|_{2}^{2}.$
@@ 4.3.3. Prototype Learning Regularization
(9) $\mathcal{L}_{div}=\frac{1}{M}\sum_{i=1}^{M}\sum_{j=i+1}^{M}\max(0,sim(P_{i},P_{j})-\theta_{min}),$
@@
(10) $\mathcal{L}=\mathcal{L}_{pred}+\lambda_{con}\mathcal{L}_{con}+\lambda_{rep}\mathcal{L}_{rep}+\lambda_{div}\mathcal{L}_{div}.$
@@
8:         Inference all the extracted subsets $\mathcal{S}$ using $\Phi$ on $\mathcal{D}_{S}$;
@@
Figure 4. Training procedure of LGDESetNet.
4.3.4. Embedding-Projection Training Procedure
Training discrete prototypes directly using gradient descent on randomly initialized values is challenging due to the non-contiguous nature of prototype representation. Direct L1 regularization could incur stringent constraints and adversely affect the model performance due to the skill sparsity. To address these limitations, we introduce an embedding-projection iterative learning algorithm for discrete prototypes. Algorithm 1 details the training procedure with Figure 4 graphically illustrating the workflow. The algorithm proceeds in five steps: (1) We start by optimizing the target embeddings $E^{\text{p}}$ for prototypes, updating them without back-propagating gradients to the discrete representations. Since $E^{\text{p}}$ is contiguous, gradient descent can effectively optimize it. This stage includes $N_{w}$ warm-up epochs to guide $E^{\text{p}}$ towards the appropriate semantic space. (2) Once retrieving the continuous prototype embeddings from the trained model, (3) we identify the closest matching subsets from frequent skill sets in the training data, (4) and then update the discrete prototypes and their corresponding embeddings $E^{\text{p}*}$. (5) Upon finalizing, we enhance the prototype’s adaptability by adding a minimal bias term to its discrete representation. Specifically, we freeze $E^{\text{p}*}$ and introduce a learnable bias term $\triangle E^{\text{p}}$, whose weight vector $\triangle P$ is initialized as zeros and L1-regularized. This fine-tuning step preserves explainability while allowing for nuanced expression of prototypical semantic uncertainty. To demonstrate the effectiveness of our architecture design for set-based modeling, we provide a theoretical proof of network permutation invariance in the Appendix A.
4.4. Complexity Analysis
We present a theoretical analysis of LGDESetNet’s computational complexity to demonstrate its feasibility for real-world deployment. The model’s time complexity is dominated by two primary components: the disentangled discrete subset selection module with complexity $O(n^{2}\times d\times H)$ for input sets of size $n$ with embedding dimension $d$ across $H$ views, and the prototypical set learning component requiring $O(H\times M\times d)$ operations for comparing extracted subsets with $M$ prototypes. Combined, the overall inference time complexity is $O(n^{2}\times d\times H+H\times M\times d)$, which remains manageable for typical skill set sizes. The quadratic dependency on input set size $n$ aligns with attention-based methods in set processing, including Set-Tree (Hirsch and Gilad-Bachrach, 2021) and SESM (Geng et al., 2020), and remains computationally feasible given the typically modest number of skills in real-world job postings. Importantly, our model’s linear scaling with respect to views ($H$) and prototypes ($M$) enables precise control over computational requirements and model capacity, allowing for effective deployment across various computational environments. For example, reducing $H$ decreases skill set selection costs, while limiting $M$ reduces prototype comparison operations. The skill/set-wise information fusion modules can be simplified to constant-time operations when external information is unnecessary, further optimizing computational requirements. These properties ensure LGDESetNet remains computationally efficient while preserving explainability across diverse applications.
Table 2. Detailed Dataset Features
@@ 4.4. Complexity Analysis
5. Experiments
Table 3. Overall performance evaluation. The listed methods include two categories: salary prediction models and self-explainable models. The best results are highlighted in bold.
Methods Skill-Salary Job Postings (Sun et al., 2021) Job Salary Benchmarking (Meng et al., 2022a)
@@ 5. Experiments
5.1. Experimental Setup
5.1.1. Datasets
The data descriptions are provided in Section 3. We use four real-world, publicly available job posting datasets with distinct job and skill distributions: (1) Skill-Salary Job Postings datasets (Sun et al., 2023), including IT and designer topics. (2) Job Salary Benchmarking datasets (Meng et al., 2022b), covering high-tech and financial industries. We provide overview of detailed dataset features in Table 2, and provide more descriptive statistics in Appendix B.1.
5.1.2. Baseline Methods
@@
  • HSBMF (Meng et al., 2018) is a salary benchmarking model utilizing a Bayesian approach to capture the hierarchical structure of job postings.
  • SSCN (Sun et al., 2021) is an enhanced neural network with cooperative structure for separating job skills and measuring their values based on massive job postings.
  • NDP-JSB (Meng et al., 2022a) develops a non-parametric Dirichlet process-based latent factor model to jointly model the latent representations of job contexts.
We also compare LGDESetNet with the following self-explainable models for set-based modeling.
  • Set-Tree (Hirsch and Gilad-Bachrach, 2021) is a tree-based general framework for processing sets and quantifies the relative importance of each item according to its frequency.
  • SESM (Geng et al., 2020) is an instance-wise selection method that discovers a fixed number of sub-inputs to explain its own prediction process.
  • ProtoPNet (Chen et al., 2019) is the pioneering work in prototype-based models, which dissects the input for more fine-grained interpretation.
  • TesNet (Wang et al., 2021) is another prototype-based image recognition deep network that introduces a transparent embedding space for disentangled global prototype construction.
  • ProtoConcepts (Ma et al., 2024) modifies prototype geometry to enable visualizations of prototypes from multiple training images.
@@ 5.1.3. Implementation Details
We evaluate the performance based on RMSE/MAE for salary prediction tasks. All results are reported based on 5 rounds of experiments. The datasets are partitioned into training/validation/testing sets by a ratio of 6:2:2. For fair comparison, we select the hyperparameter configurations based on the best performance on the validation set. For all experiments, prototype number $M$ is set to 64 and head number $H$ is set to 4. The loss function hyperparameters $\lambda_{rep}$, $\lambda_{con}$, and $\lambda_{div}$ are all set to 0.1. We train our model with ADAM optimizer (Kingma and Ba, 2014). The transformation layer $\mathcal{T}$ is implemented using a 2-layer MLP layer with 256 hidden units. Experiments were implemented in Pytorch and run on a server with 8 Nvidia GeForce RTX 3090 GPUs.  [^1] Experiment code is available at https://anonymous.4open.science/r/TKDD-Submission-LGDESetNet.
5.2. Overall Performance Evaluation

### v3

@@ 4.2.2 Skill-wise External Information Fusion
$\begin{split}E^{\text{s}}_{h}&=\text{pool}\left(\{\alpha^{\text{lv}}_{i}\phi(x_{i})|x_{i}\in S_{h}\}\right).\end{split}$ (3)
@@
Figure 3: Illustrative examples of co-occurrence graphs. Low-weight edges are eliminated to ease presentation.
4.2.3 Co-occurrence Graph Density Regularization
To enhance explainability and robustness, extracted skill subsets should reflect meaningful real-world interactions. Yet, focusing solely on prediction performance could generate subsets cluttered with irrelevant skills, reducing semantic clarity. To address this, we introduce a regularization approach aimed at refining these subsets to meet key intuitive criteria: (i) Semantic coherence within subsets, (ii) Distinct semantic boundaries between different subsets, and (iii) Avoidance of trivial subset selections (neither fully inclusive nor empty skill sets).
Observing that skills frequently appeared together in job postings likely share semantic connections (e.g., “Machine Learning”, “Algorithms”, and “Python” often cluster in AI roles), we leverage global semantic interactions inferred from such co-occurrence patterns. Direct co-occurrence analysis, however, is computationally expensive and hampered by data sparsity. Therefore, we model these relationships using a skill co-occurrence graph $G_{skill}=\langle U,L\rangle$, drawing from skill pair associations across the dataset, where $U$ is the universal set of skills and $L$ represents the links indicating co-occurrence frequencies between skill pairs. As exemplified in Figure 3, each graph node corresponds to a skill, and a link $l_{i,j}$ between nodes $x_{i}$ and $x_{j}$ signifies their connectivity, quantified by their co-occurrence frequency $w_{l_{i,j}}$. To quantify semantic connections within subsets, we design a graph density-based [60] regularization, focusing on the subgraph $G_{h}=\langle S_{h},L_{G_{h}}\rangle$ for each subset $S_{h}$. The subgraph’s density, calculated as $\frac{\sum_{l\in L_{\mathrm{G}_{h}}}w_{l}}{|L_{{\mathrm{G}_{h}}|}}$, gauges the semantic cohesiveness of the subset. By maximizing this density, we encourage subsets that are semantically well-connected, while minimizing trivial or overly inclusive selections. Our graph regularization of the extracted subsets $\{S_{h}\}_{h=1}^{H}$ is thus defined as:
$\mathcal{L}_{con}=-\log\frac{1}{H}\sum_{h=1}^{H}\frac{\sum_{l\in L_{\mathrm{G}_{h}}}w_{l}}{|L_{\mathrm{G}_{h}}|}.$ (4)
@@ 4.2.3 Co-occurrence Graph Density Regularization
4.3 Prototypical Set Learning
Based on extracted subsets, LGDESetNet matches the input with learnable prototypes representing global semantics and accordingly models the output as an aggregation of the global contributions of these prototypes.
4.3.1 Set-oriented Prototypical Part Layer
@@
We assess the alignment between the input set and each prototype (i.e., global semantics) by comparing the similarity of extracted subsets to the prototypes. Notably, the discrete set representation brings the curse of dimensionality, complicating set similarity estimation. For instance, “C” and “C++” are treated as distinct skills despite their semantic closeness, potentially hindering model convergence and accuracy. To address this, we project prototype $P_{i}$ and subset $S_{h}$ into semantic representation vectors with the same skill embedding table as the input set, denoted as $E^{\text{p}}_{i}=\text{pool}\left(\{\gamma^{\text{s}}_{i}\gamma^{\text{lv}}_{i}\phi(x_{i})|x_{i}\in U\}\right)$ and $E^{\text{s}}_{h}=\text{pool}\left(\{\alpha^{\text{lv}}_{i}\phi(x_{i})|x_{i}\in S_{h}\}\right)$. Considering the distribution shift between the whole set and single-view subsets, we introduce a transformation layer $\mathcal{T}$ to further project the subsets and prototypes into a subset representation space as $Z^{\text{p}}_{i}=\mathcal{T}\left(E^{\text{p}}_{i}\right)$ and $Z^{\text{s}}_{h}=\mathcal{T}\left(E^{\text{s}}_{h}\right)$. In this way, we better capture semantic-level associations while ensuring the similarity can be measured by aligned representation of prototypes and extracted subsets.
@@
$sim(S_{h},P_{i})=\text{log}\left(\frac{\|Z^{\text{s}}_{h}-Z^{\text{p}}_{i}\|_{2}^{2}+1}{\|Z^{\text{s}}_{h}-Z^{\text{p}}_{i}\|_{2}^{2}+\epsilon}\right).$ (5)
@@
$y=\sum_{i=1}^{M}\gamma^{\text{sal}}_{i}\cdot\text{softmax}(\sum_{h=1}^{H}sim(S_{h},P_{i})).$ (6)
4.3.2 Set-wise External Information Fusion
In practice, each skill set is enriched with set-wise external information, which is also useful in determining salary outcomes. For example, job salaries are highly correlated with the work experience and the pricing level of the base city. However, directly fusing such information into deep and wide embeddings [61] could obscure the model’s explainability. Therefore, we separate the set-wise external information from skill embeddings, acknowledging that the salary effects of global prototypical skill sets may vary across different job scenarios. Specifically, we model this prototype salary effect dynamic as a function of the set-wise external features $C=\{c_{1},c_{2},...,c_{l}\}$. Drawing inspiration from DeepFM [62], we utilize a factorization machine encoder for low-order feature interactions and an MLP encoder for high-order interactions. By aggregating the embeddings from both encoders, we then employ another MLP to estimate the salary effects of each prototype, ensuring a dynamic and contextual assessment of skill sets’ impact on salary predictions:
$\begin{split}\gamma^{\text{sal}}&=\text{MLP}\left(\text{FM}(C)+\text{MLP}(c_{1}|c_{2}|...|c_{l})\right).\\ \text{FM}(C)&=w_{0}+\sum_{i=1}^{l}w_{i}c_{i}+\sum_{i=1}^{l}\sum_{j=i+1}^{l}c_{i}\odot c_{j},\end{split}$ (7)
@@ 4.3.2 Set-wise External Information Fusion
4.3.3 Prototype Learning Regularization
To enhance explainability, we introduce regularization methods for our prototypical set learning procedure, whose basic ideas are widely adopted in prototype learning for other data structures [50, 47]. First, since prototypes are desired to be representative exemplars for the subsets, we introduce a representation regularization term to encourage each subset to be close to at least one prototype and encourage each prototype to be close to at least one extracted subset.
$\mathcal{L}_{rep}=\frac{1}{H}\sum\limits_{h=1}^{H}\min_{i\in[1,M]}\|Z^{\text{s}}_{h}-Z^{\text{p}}_{i}\|_{2}^{2}+\frac{1}{M}\sum_{i=1}^{M}\min_{h\in[1,H]}\|Z^{\text{s}}_{h}-Z^{\text{p}}_{i}\|_{2}^{2}.$ (8)
@@ 4.3.3 Prototype Learning Regularization
$\mathcal{L}_{div}=\frac{1}{M}\sum_{i=1}^{M}\sum_{j=i+1}^{M}\max(0,sim(P_{i},P_{j})-\theta_{min}),$ (9)
@@
$\mathcal{L}=\mathcal{L}_{pred}+\lambda_{con}\mathcal{L}_{con}+\lambda_{rep}\mathcal{L}_{rep}+\lambda_{div}\mathcal{L}_{div}.$ (10)
@@
8:         Infer all the extracted subsets $\mathcal{S}$ using $\Phi$ on $\mathcal{D}_{S}$;
@@
Figure 4: Training procedure of LGDESetNet.
4.3.4 Embedding-Projection Training Procedure
Training discrete prototypes directly using gradient descent on randomly initialized values is challenging due to the non-contiguous nature of prototype representation. Direct L1 regularization could incur stringent constraints and adversely affect the model performance due to the skill sparsity. To address these limitations, we introduce an embedding-projection iterative learning algorithm for discrete prototypes. Algorithm 1 details the training procedure with Figure 4 graphically illustrating the workflow. The algorithm proceeds in five steps: (1) We start by optimizing the target embeddings $E^{\text{p}}$ for prototypes, updating them without back-propagating gradients to the discrete representations. Since $E^{\text{p}}$ is contiguous, gradient descent can effectively optimize it. This stage includes $N_{w}$ warm-up epochs to guide $E^{\text{p}}$ towards the appropriate semantic space. (2) Once retrieving the continuous prototype embeddings from the trained model, (3) we identify the closest matching subsets from frequent skill sets in the training data, (4) and then update the discrete prototypes and their corresponding embeddings $E^{\text{p}*}$. (5) Upon finalizing, we enhance the prototype’s adaptability by adding a minimal bias term to its discrete representation. Specifically, we freeze $E^{\text{p}*}$ and introduce a learnable bias term $\triangle E^{\text{p}}$, whose weight vector $\triangle P$ is initialized as zeros and L1-regularized. This fine-tuning step preserves explainability while allowing for nuanced expression of prototypical semantic uncertainty. To demonstrate the effectiveness of our architecture design for set-based modeling, we provide a theoretical proof of network permutation invariance in Appendix A.
Table 2: Detailed Dataset Features
@@ 4.3.4 Embedding-Projection Training Procedure
4.4 Complexity Analysis
We present a theoretical analysis of LGDESetNet’s computational complexity to demonstrate its feasibility for real-world deployment. The model’s time complexity is dominated by two primary components: the disentangled discrete subset selection module with complexity $O(n^{2}\times d\times H)$ for input sets of size $n$ with embedding dimension $d$ across $H$ views, and the prototypical set learning component requiring $O(H\times M\times d)$ operations for comparing extracted subsets with $M$ prototypes. Combined, the overall inference time complexity is $O(n^{2}\times d\times H+H\times M\times d)$, which remains manageable for typical skill set sizes. The quadratic dependency on input set size $n$ aligns with attention-based methods in set processing, including Set-Tree [23] and SESM [59], and remains computationally feasible given the typically modest number of skills in real-world job postings. Importantly, our model’s linear scaling with respect to views ($H$) and prototypes ($M$) enables precise control over computational requirements and model capacity, allowing for effective deployment across various computational environments. For example, reducing $H$ decreases skill set selection costs, while limiting $M$ reduces prototype comparison operations. The skill/set-wise information fusion modules can be simplified to constant-time operations when external information is unnecessary, further optimizing computational requirements. These properties ensure LGDESetNet remains computationally efficient while preserving explainability across diverse applications.
5 Experiments
Table 3: Overall performance evaluation. The listed methods include two categories: salary prediction models and self-explainable models. The best results are highlighted in bold.
Methods Skill-Salary Job Postings [4] Job Salary Benchmarking [8]
@@ 5 Experiments
5.1 Experimental Setup
5.1.1 Datasets
The data descriptions are provided in Section 3. We use four real-world, publicly available job posting datasets with distinct job and skill distributions: (1) Skill-Salary Job Postings datasets [55], including IT and designer topics. (2) Job Salary Benchmarking datasets [56], covering high-tech and financial industries. We provide overview of detailed dataset features in Table 2, and provide more descriptive statistics in Appendix B.1.
5.1.2 Baseline Methods
@@
  • HSBMF [7] is a salary benchmarking model utilizing a Bayesian approach to capture the hierarchical structure of massive job postings.
  • SSCN [4] is an enhanced neural network with cooperative structure for separating job skills and measuring their values based on job postings.
  • NDP-JSB [8] develops a non-parametric Dirichlet process-based latent factor model to jointly model the latent representations of job contexts.
We also compare LGDESetNet with the following self-explainable models for set modeling.
  • Set-Tree [23] is a tree-based general framework for processing sets and quantifies the relative importance of each item according to its frequency.
  • SESM [59] is an instance-wise selection method that discovers a fixed number of sub-inputs to explain its own prediction.
  • ProtoPNet [14] is the pioneering work in prototype-based models, which dissects the input images into patches for fine-grained interpretation.
  • TesNet [44] is another prototype-based image recognition deep network that introduces a transparent embedding space for disentangled global prototype construction.
  • ProtoConcepts [43] modifies prototype geometry to enable visualizations of prototypes from multiple training images.
@@ 5.1.3 Implementation Details
We evaluate the performance based on RMSE/MAE for salary prediction tasks. All results are reported based on 5 rounds of experiments. The datasets are partitioned into training/validation/testing sets by a ratio of 6:2:2. For fair comparison, we select the hyperparameter configurations based on the best performance on the validation set. For all experiments, prototype number $M$ is set to 64 and head number $H$ is set to 4. The loss function hyperparameters $\lambda_{rep}$, $\lambda_{con}$, and $\lambda_{div}$ are all set to 0.1. We train our model with ADAM optimizer [63]. The transformation layer $\mathcal{T}$ is implemented using a 2-layer MLP layer with 256 hidden units. Experiments were implemented in Pytorch and run on a server with 8 Nvidia GeForce RTX 3090 GPUs. 111 Code is available at https://anonymous.4open.science/r/FCS-Submission-LGDESetNet.
Figure 5: Ablation studies across four datasets.
5.2 Overall Performance Evaluation